# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v1

@@ Methods
Datasets. We use the most comprehensive benchmark dataset, HLSyn (Bai et al. 2023). It consists of 42 kernels covering various categories: linear algebra on vectors and matrices, data mining, stencil operations, etc. We utilize the AMD/Xilinx HLS tool, Vitis 2021.1 (AMD/Xilinx 2020), to run HLS targeting the Xilinx Alveo U200 FPGA with a working frequency of 250MHz. We select 6 kernels as the target kernels that span representative categories including linear algebra computation, data mining, and stencil computation, and we use the other kernels as source kernels. Table 2 shows the dataset statistics. We introduce the kernels’ details in the appendix. The HLSyn dataset was generated by running the bottleneck-based heuristics method, AutoDSE (Sohrabizadeh et al. 2021), for 24 hours. Many designs explored by AutoDSE are unavailable in HLS, so we collect this information to train a HARP classification model. Since its accuracy already exceeds 95%, there is no need to employ MoE on the classification model. We use the available designs in the dataset to train the regression model, and we employ MoE on the regression model.
@@
#Class 36185 418 428 526 1421 1837 1651
#Regre 9086 77 231 121 348 257 103
Models. Based on our pre-exploration, we use 4 experts in the low-level MoEs, and we set the weight of the regularization term of both low-level and high-level MoE, $\alpha$ and $\beta$, to 5e-3. Our baselines include the SOTA GNN method, HARP (Sohrabizadeh et al. 2023); HARP+MAML (Bai et al. 2022), which applies the meta-learning method MAML (Finn, Abbeel, and Levine 2017) on this task proposed by (Bai et al. 2022), since MAML could learn a more generalizable initialization of parameters; and ProgSG (Qin et al. 2024), which combines HARP and language model.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. As introduced in the preliminary, we conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned model’s mean squared error (MSE) on the left-out data points in the target kernels. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs and the time of an hour, since it typically takes about an hour to traverse 75,000 designs. We use the fine-tuned classification model to predict the validity and the fine-tuned regression model to predict the FPGA’s latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies. We run HLS for them and pick up the best valid design that has the lowest latency. Since we have used the dataset to fine-tune, we pick the best design from the selected top-10 designs and the dataset, and calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then report the mean results.
@@
Table 1 shows the main results. In the main experiments, we use the second design of the high-level gating network as it performs better. Neither MAML nor ProgSG performs well in our setting. It might be because we have many more source kernels compared to the previous paper that utilizes MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable meta-learning. ProgSG which combines GNN and language models has a strong ability on the HLSyn dataset when the training data is sufficient (Qin et al. 2024), but since the language model needs a large dataset to fine-tune, it severely overfits in our data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the hierarchical MoE is a more generalizable structure. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP, but different kernels favor different low-level MoEs. For example, the block MoE model performs better than the node MoE and the graph MoE for “Fd”, while graph MoE is the best for “Gemv”, “Ja”, and “Tr”. Since a single MoE model could not perform well on all kernels, its overall MSE and the average speedup might not outperform HARP. By aggregating them together, the hierarchical MoE performs the best or close to the best on almost every kernel. Its geometric mean speedup achieves about 26.6% higher than HARP.
@@
Various model structures. To verify the effectiveness of our elaborate model design, we run extensive experiments on various model structures. First, instead of aggregating the three low-level MoEs, can we apply MoE on the three granularities together in a single model? We specify the detailed structure design in the appendix. Table 3 shows the results. When we use MoE on two or three granularities in a single model, the loss is usually higher than the lowest loss when using only one granularity. Apart from selecting 50 representative data points by K-means for fine-tuning, we also experiment with random data split. We use the same random split for all models to ensure fair comparison. Hierarchical MoE is the most stable model when it faces low-quality fine-tuning data in the random split. If we stack MoE on various granularities in the same model, the structure might be too complex and thus unstable to train on a small dataset.
@@
Second, does the hierarchical MoE benefit from combining the three granularities? To answer it, we still use the hierarchical MoE structure, but we use the same low-level MoE model as the three experts of the high-level MoE. Table 4 shows the results. It verifies that aggregating the three granularities performs the best. Using only the node MoE or graph MoE also performs well in the online evaluation, and this might be due to the increased number of experts. Nonetheless, aggregating the three granularities could further improve the performance. Third, to verify whether the performance gain simply comes from having more parameters, we enlarge the hidden size of the HARP model from 64 to 128 or 256, and the model loss increases. Due to the space limit, the results are in the appendix. Therefore, the key is the MoE structure instead of a larger parameter size.
@@
Fourth, we conduct an ablation study on the two-stage training and the high-level gating network’s design, and the results are listed in Table 5. There are two components in the two-stage training: the warmup epochs, and training three experts jointly and separately in turn after the warmup. We either disable the alternative training or the warmup. The lowest MSE is achieved when we use both of them. Also, the high-level gating network based on hidden representations is better than that based on input features.
@@
We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes but not normal nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 7. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves domain generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. Based on the unique challenges and opportunities of this task, we propose the hierarchical MoE structure, where different expert networks can specialize in different regions, and it can be flexibly adapted to any GNN model. In the low-level MoE, we apply MoE on the three natural granularities of the graph: nodes, basic blocks, and the graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly decide which granularity to utilize MoE. To address the severe expert polarization, we propose two-stage training via deactivating the high-level gating network for some epochs. Extensive experiments have verified the effectiveness of hierarchical MoE. We focus on the generalizability of the HLS prediction task, but our model might also potentially improve the generalizability in other applications of various domains.
@@ Acknowledgments
This work was partially supported by NSF grants 2211557, 1937599, 2119643, and 2303037, SRC JUMP 2.0 PRISM Center, NASA, Okawa Foundation, Amazon Research, Cisco, Picsart, Snapchat, and the CDSC industrial partners (https://cdsc.ucla.edu/partners/). The authors would also like to thank AMD/Xilinx for HACC equipment donation, and Marci Baun for editing the paper. J. Cong has a financial interest in AMD.
@@ Appendix A Additional Background Knowledge of HLS
Table 8: Introduction of pragmas.
Pragma type Option Meaning
Parallel An integer Number of loop iterations to parallelize
Pipeline ”Off” No pipeline
”Flatten” Fine-grained pipeline
NA Coarse-grained pipeline
Loop tiling An integer The factor of loop tiling
There are three types of pragmas: parallel, pipeline, and loop tiling. For the parallel pragma placeholders, we should insert an integer as the parallel factor. For the pipeline pragma placeholders, we should choose between the fine-grained pipeline, coarse-grained pipeline, or not using the pipeline. For the loop tiling pragma placeholders, we should insert an integer as the loop tiling factor. Table 8 lists the introduction of the pragmas.
Designing the pragmas is labor-intensive and heavily relies on hardware knowledge, which motivates the automated pragma design research. Different pragmas can lead to significantly different FPGA performance. Table 9 shows the latency (clock cycle) of each target kernel’s best pragma design and worst pragma design in our dataset. The shorter latency the better. We also list the mean and standard deviation of latency in our dataset, and the latency of not inserting any pragma. We can see that the gap between the best design and worst design is very big, and the worst design is sometimes even worse than not inserting any pragma. This is because some pragmas might use too many resources and occupy the resources of other modules. The standard deviation is also very large, and sometimes it is even larger than the mean. Therefore, the pragma design is very difficult.
Table 9: The latency (clock cycle) of each target kernel’s best pragma design and worst pragma design in our dataset. We also list the mean and standard deviation of latency in our dataset, and the latency of not inserting any pragma.
@@ Appendix B Additional Problem Analysis
Domain shift brings a lot of challenges to our task. First, as shown in Table 9, the mean latency of different kernels is very different, and we hope the model can learn this distribution, so we cannot align the representation space of different kernels, as did in many previous domain generalization papers. We extract the graph representation of each design point from the HARP model and visualize it in Figure 4. In the figures, points belonging to the same kernel have the same color. Due to the limited number of colors, different kernels might use the same color. We can see that points from the same kernels are clustered together, while points from different kernels are located in different places. The visualization further verifies that the domain difference between different kernels is very big. Such a big difference is informative to the model, so we cannot align the representation space of different kernels.
@@
Second, kernels share some similar substructions, where both their similarities and differences should be carefully noticed. The similarity could benefit the MoE model by utilizing the knowledge learned from other kernels. Nonetheless, the difference is also very important. An example is the “syr2k” kernel and the “syrk” kernel. They only differ in one line inside a nested loop, shown in Figure 5 and Figure 3. However, their best pragma designs are very different. In the best pragma design of “syrk”, the tiling factor of “__TILE__L2” is 60, but it is 1 for “syr2k”; the parallel factor of “__PARA__L1” and “__PARA__L2” are 20 and 1 for “syrk”, but they are 8 and 4 for “syr2k”; “syrk” uses coarse-grained pipeline for “__PIPE__L2”, but “syr2k” does not use pipeline there. Therefore, domain generalization is difficult for HLS prediction models. The MoE structure can utilize similar experts to deal with similarities, but it can also discern the differences and use different experts to deal with them.
@@
Figure 6 illustrates the HARP model using the same diagram style as our model’s diagram (Figre 2 in the main paper). It shows the five components of the HARP model. Since we focus more on the MoE structure, the diagram ignores some details of the HARP model. For more details, please refer to the HARP paper (Sohrabizadeh et al. 2023).
@@ Appendix C Detailed Experiment Settings
Datasets. The HLSyn benchmark dataset (Bai et al. 2023) consists of 42 kernels selected from the MachSuite dataset (Reagen et al. 2014) and the Polyhedral dataset (PolyBench) (Pouchet 2012). Since the FPGA’s latency is too diverse, it is difficult for the model to directly learn. Thus, we normalize the FPGA’s latency by $0.5*log_{2}(\frac{1e7}{\text{latency}})$ as the normalized performance. It is the lower the better for the original latency, but it is the higher the better for the normalized performance. Our HLS targets the Xilinx Alveo U200 FPGA with a working frequency of 250MHz. We also normalize the resource utilization by dividing it by the total available resources of the targeting FPGA. We mainly care about four types of resources: LUT, FF, DSP, and BRAM, so we have five prediction targets including the normalized performance.
@@
Our resource constraint is that each resource type should be used no more than 80%, because we need to leave some spaces for downstream tasks or other additional modules. We hope the model can make accurate predictions of the performance and resource utilization, so we can search for the best pragma design that satisfies the resource constraints and has the best normalized performance (lowest latency). We rely on the classification model to help us wipe out the invalid designs, but we rely on the regression model to help us wipe out the valid designs that exceed our resource constraints.
@@
Evaluation. We use 80% of data points in the source kernels as the training set to pre-train the model, 10% for validation, and 10% for testing. We use five random seeds (1, 2, 3, 4, 5) to randomly split the source kernel’s dataset. The MSE on the source kernels is unimportant, so currently the test set on the training kernels is not used. We use the validation set for early stopping. We pretrain the model for 1000 epochs, then we finetune it for 500 epochs on the target kernels. During fine-tuning, we use the same random seed as used in that model’s pretraining. We run offline evaluation five times. For online evaluation, since HLS takes a long time, we use the three models trained with random seeds 1, 2, and 3 to run HLS. In offline evaluation, we calculate the sum of the MSE of the five prediction targets. In online evaluation, from the designs that are predicted to be valid and under resource constraints, we pick the top 10 designs with the lowest predicted latencies to run HLS. When running HLS, we set the timeout as 3 hours. If the HLS does not finish within 3 hours, it is very likely to be unsynthesizable, so we will terminate the HLS and treat it as an invalid design. We calculate the mean speedup of the three repeated HLS experiments for each target kernel. Then we calculate the mean speedup or geometric mean speedup for all target kernels.
@@
Baseline methods. For all baseline methods, we follow their original best hyper-parameter settings. For our models, we follow the same hyper-parameter setting and training strategy as HARP, except for the MoE’s unique hyper-parameters. We will specify the hyper-parameters in a later section.
  • HARP: we use its original settings and official codes (https://github.com/UCLA-VAST/HARP).
  • HARP+MAML: we use MAML to help find a good initialization for HARP model’s parameters based on the source kernels. Then we use the normal fine-tuning strategy for it on the target kernels. We use the implementation of the “learn2learn” Python package. We randomly sample 32 kernels with replacement per epoch to train.
  • ProgSG: it uses both text embedding of the source code and the ProGraML graph as the input. We use its official codes (https://github.com/zongyueqin/progsg).

### v2

@@ Methods
Datasets. We use one of the most comprehensive benchmark datasets, HLSyn (Bai et al. 2023). It contains 42 kernels covering various categories: linear algebra of vectors and matrices, data mining, stencil operations, etc. We utilize the AMD/Xilinx HLS tool, Vitis 2021.1 (AMD/Xilinx 2020), to run HLS targeting the Xilinx Alveo U200 FPGA with a working frequency of 250MHz. We select 6 kernels as the target kernels that span representative categories including linear algebra, data mining, and stencil. The other kernels are source kernels. Table 2 shows the dataset statistics. We introduce their details in the appendix. The HLSyn dataset was generated by running the heuristics method, AutoDSE (Sohrabizadeh et al. 2021), for 24 hours. Many designs explored by AutoDSE are unavailable in HLS, so we collect this information to train a HARP classification model. Since its accuracy already exceeds 95%, there is no need to employ MoE on the classification model. We use the available designs in the dataset to train the regression model, and we employ MoE on the regression model.
@@
#Class 37697 418 428 526 1421 1837 1651
#Regre 9418 77 231 121 348 257 103
Models. Based on our pre-explorations, we use 4 experts in the low-level MoEs, and we set the regularization terms’ weights of both low-level and high-level MoEs to 5e-3. Our baselines include the SOTA GNN model, HARP (Sohrabizadeh et al. 2023); HARP+MAML (Bai et al. 2022), which applies meta-learning method MAML (Finn, Abbeel, and Levine 2017) on this task inspired by (Bai et al. 2022) to learn a more generalizable initialization of parameters; and ProgSG (Qin et al. 2024), which combines HARP and language model. Although hierarchical MoE brings many more parameters, it needs about 32 GB of memory and can still fit into a single GPU. HARP and our model contain 359,370 and 1,329,403 parameters respectively.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. We conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned regression model’s mean squared error (MSE) on the left-out data points in the target kernels for each regression objective and sum up the five objectives’ MSE. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs, since it typically takes an hour. We use the fine-tuned classification model to predict the validity and the fine-tuned regression model to predict the latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies to run HLS. We report the best design from the selected top-10 and the training dataset of the target kernels during fine-tune, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then we report the mean results.
@@
Table 1 shows the main results. In the main experiments, we use the second design of the high-level gating network as it performs better. MAML could not perform well in this setting. It might be because we have many more source kernels compared to the previous paper that utilizes MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable meta-learning. ProgSG which combines GNN and language models has a strong ability on the HLSyn dataset when the training data is sufficient (Qin et al. 2024), but it severely overfits in our data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the hierarchical MoE is a more generalizable structure. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP. However, different kernels favor different low-level MoEs. For example, the block MoE model performs better than node MoE and graph MoE for “Fd”, while graph MoE is the best for “Gemv”, “Ja”, and “Tr”. By aggregating them together, the hierarchical MoE performs the best or close to the best on almost every kernel. Its geometric mean speedup achieves about 26.6% higher than HARP.
Table 3: Offline evaluation results on more complex kernels.
HARP HARP+MAML Hierarchical MoE
MSE 0.426 3.227 0.401
Complex kernels. Kernels in the HLSyn dataset are generally small. We further test the models on 5 more complex kernels, including ”3d-rendering” and ”spam-filter” from the Rosetta benchmark (Zhou et al. 2018) and three self-constructed kernels, which are introduced in the appendix. They are closer to real-world applications. Following previous settings, we train the regression model on all HLSyn kernels and finetune it on 50 points per kernel on the new kernels. The results are shown in Table 3. It is challenging for all models to perform well, while the hierarchical MoE model performs the best. It remains a future work to generally enhance model capacity on complex kernels.
Ablation Study
@@
To verify the effectiveness of our model design, we run extensive experiments on various model structures as ablation studies. First, instead of aggregating the three low-level MoEs, can we apply MoE on the three granularities together in a single model? We specify the detailed structure design in the appendix. Table 4 shows the results. When we use MoE on two or three granularities in a single model, the loss is usually higher than the lowest loss when only using MoE on one granularity. Apart from selecting 50 representative data points by K-means for fine-tuning, we also experiment with random data split. We use the same random split for all models to ensure fair comparison. Hierarchical MoE is the most stable model when it faces low-quality fine-tuning data in the random split. If we stack MoE on various granularities in the same model, the structure might be too complex and thus unstable to train on a small dataset.
@@
Second, does the hierarchical MoE benefit from combining the three granularities? To answer it, we still use the hierarchical MoE structure, but we use the same low-level MoE model as the three experts of the high-level MoE. Table 5 shows the results. It verifies that aggregating the three granularities performs the best. Using only the node MoE or graph MoE also performs well in the online evaluation, and this might be due to the increased number of experts. Nonetheless, aggregating the three granularities could further improve the performance.
@@
Third, we conduct an ablation study on the two-stage training and the high-level gating network’s design, and the results are listed in Table 6. There are two components in the two-stage training: the warmup epochs, and training three experts jointly and separately in turn after the warmup. We either disable the alternative training or the warmup. The lowest MSE is achieved when we use both of them. Also, the high-level gating network based on hidden representations is better than that based on input features.
There are more ablation studies in the appendix. They verify that the hierarchical MoE’s performance is not due to increased parameter size or increased expert number, but the hierarchical structure and combination of three granularities. An ablation study also verifies that the current complicated design of the block MoE is better than some other variants.
@@
We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 8. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves the generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. Based on the unique challenges and opportunities, we propose the hierarchical MoE structure, where different expert networks can specialize in different regions, and it can be flexibly adapted to any GNN. In the low-level MoE, we apply MoE on the three natural granularities of the graph: node, basic block, and graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly decide which one to use. To address the severe expert polarization, we propose a two-stage training strategy. Extensive experiments have verified its effectiveness. Nonetheless, the generalizability of HLS prediction models still remains a big challenge, and future works can further improve the performance.
@@ Acknowledgments
This work was partially supported by NSF grants 2211557, 1937599, 2119643, 2303037, and 2312501, SRC JUMP 2.0 PRISM Center, Amazon Research, Snapchat, and the CDSC industrial partners (https://cdsc.ucla.edu/partners/). The authors would also like to thank AMD/Xilinx for HACC equipment donation, and Marci Baun for editing the paper. J. Cong has a financial interest in AMD.
@@ Appendix A Additional Background Knowledge of HLS
Following HARP (Sohrabizadeh et al. 2023), we develop the model based on the open-source AMD/Xilinx Merlin Compiler (Cong et al. 2016), because it reduces the complex HLS pragmas to only three Merlin pragmas, greatly benefiting the pragma search process. It applies source-level code transformations to enable various automatic architectural optimizations such as memory burst, memory coalescing, and coarse-grained optimizations. There are three types of Merlin pragmas: parallel, pipeline, and loop tiling. For the parallel pragma placeholders, we should insert an integer as the parallel factor. For the pipeline pragma placeholders, we should choose between the fine-grained pipeline, coarse-grained pipeline, or not using the pipeline. For the loop tiling pragma placeholders, we should insert an integer as the loop tiling factor.
Designing the pragmas is labor-intensive and heavily relies on hardware knowledge, which motivates the automated pragma design research. Different pragmas can lead to significantly different FPGA performance. Table 9 shows the latency (clock cycle) of each target kernel’s best pragma design and worst pragma design in the HLSyn dataset. The shorter latency the better. We also list the mean and standard deviation of latency in the dataset, and the latency of not inserting any pragma. We can see that the gap between the best design and worst design is very big, and the worst design is sometimes even worse than not inserting any pragma. This is because some pragmas might use too many resources and occupy the resources of other modules. The standard deviation is also very large, and sometimes it is even larger than the mean. Therefore, choosing a good pragma design is very important to the FPGA’s performance.
Table 9: The latency (clock cycle) of each target kernel’s best pragma design and worst pragma design in the HLSyn dataset. We also list the mean and standard deviation of latency in the HLSyn dataset, and the latency of not inserting any pragma.
@@ Appendix B Additional Problem Analysis
Domain shift brings a lot of challenges to our task. First, different kernels have large differences. As shown in Table 9, the mean latency of different kernels is very different. We hope the model can learn this distribution, so we cannot align the representation space of different kernels, as did in many previous domain generalization papers. We extract the graph representation of each design point from the HARP model and visualize it in Figure 4. In the figure, points belonging to the same kernel have the same color. Due to the limited number of colors, different kernels might use the same color. We can see that points from the same kernels are clustered together, while points from different kernels are located in different places. The visualization further verifies that the domain difference between different kernels is very big. Such a big difference makes the domain transfer very difficult.
@@
Second, the labeling is time-consuming on the new kernels. Each run takes minutes to hours (Sohrabizadeh et al. 2023). Therefore, we use the constraint of data scarcity of only using 50 data points per kernel to finetune the model on new kernels.
Nonetheless, a unique opportunity is that some different kernels share some similar substructions, where both their similarities and differences should be carefully noticed. The similarity could benefit the MoE model by utilizing the knowledge learned from other kernels. Nonetheless, the difference is also very important. An example is the “syr2k” kernel and the “syrk” kernel. They only differ in one line inside a nested loop, shown in Figure 5 and Figure 3. However, their best pragma designs are very different. In the best pragma design of “syrk”, the tiling factor of “__TILE__L2” is 60, but it is 1 for “syr2k”; the parallel factor of “__PARA__L1” and “__PARA__L2” are 20 and 1 for “syrk”, but they are 8 and 4 for “syr2k”. It is challenging for the model to not only benefit from the similarities and also observe the slight differences. Therefore, domain generalization is difficult for HLS prediction models. The MoE structure can utilize similar experts to deal with the similarities, but it can also discern the differences and use different experts to deal with them.
@@
Figure 6 illustrates the HARP model using the same diagram style as our hierarchical MoE model’s diagram (Figre 2). It shows the five components of the HARP model. This diagram ignores some details of the HARP model. For more details, please refer to the HARP paper (Sohrabizadeh et al. 2023).
@@ Appendix C Detailed Experiment Settings
Datasets. The HLSyn benchmark dataset (Bai et al. 2023) consists of 42 kernels selected from the MachSuite dataset (Reagen et al. 2014) and the Polyhedral dataset (PolyBench) (Pouchet 2012). Since the FPGA’s latency is too diverse, it is difficult for the model to directly learn to predict. Thus, we normalize the FPGA’s latency by $0.5*log_{2}(\frac{1e7}{\text{latency}})$ as the normalized performance. It is the lower the better for the original latency, but it is the higher the better for the normalized performance. Our HLS targets the Xilinx Alveo U200 FPGA with a working frequency of 250MHz. We also normalize the resource utilization by dividing it by the total available resources of the targeting FPGA. We mainly care about four types of resources: LUT, FF, DSP, and BRAM, so we have five prediction targets including the normalized performance.
@@
Our resource constraint is that each resource type should be used no more than 80%, because we need to leave some spaces for downstream tasks or other additional modules. We hope the model can make accurate predictions of the performance and resource utilization, so we can search for the best pragma design that satisfies the resource constraints and has the highest normalized performance (lowest latency). We rely on the classification model to help us wipe out the invalid designs, but we rely on the regression model to help us wipe out the valid designs that exceed our resource constraints.
More Complex Kernels. Since the kernels in the HLSyn benchmark dataset (Bai et al. 2023) are relatively small and simple, we also conduct experiments on more complex kernels that are closer to real-world applications. We select two kernels from the Rosetta benchmark dataset (Zhou et al. 2018): 3d-rendering and spam-filter. The “3d-rendering” kernel renders 2D images from 3D triangle mesh models. The “spam-filter” kernel trains a logistic regression model for spam email classification. The details of these two kernels can be found in the Rosetta benchmark paper.
We also construct three new kernels by ourselves. The first kernel, “att-3mm”, calculates the attention mechanism widely used in the Transformer (Vaswani et al. 2023) architecture. Its input is an embedding matrix and two parameter matrices. It uses the embedding matrix and the first parameter matrix to calculate the Q matrix, uses the embedding matrix and the second parameter matrix to calculate the K matrix, and then multiplies Q and K to get the attention matrix. The second kernel, “att-3mm-fuse”, does the same calculation but it fuses the outer loop. The third kernel, “vmmv”, is constructed by combining vector-matrix multiplication and matrix-matrix multiplication inspired by the “atax” kernel and the “gemm-p” kernel in the HLSyn benchmark.
Following the previous settings, we train the regression model on the complete HLSyn data and finetune it on 50 points per kernel on the new kernels. We use K-means to select 50 data points per kernel. We list these complex kernels’ statistics in Table 10.
Table 10: Dataset statistics of the complex kernels. “#Regre” denotes the number of regression data. The kernel acronyms represent “3d-rendering”, “spam-filter”, “att-3mm”, “att-3mm-fuse”, and “vmmv”.
HLSyn 3d Spam Att Att-fuse Vmmv
#Regre 10555 124 520 90 139 298
@@
Evaluation. During pre-training on the source kernels, we use 80% of data points as the training set, 10% for validation, and 10% for testing. We use five random seeds (1, 2, 3, 4, 5) to randomly split the source kernel’s dataset. The MSE on the source kernels is unimportant. We pretrain the model for 1000 epochs, and we select the model checkpoint with the lowest validation MSE on the validation set. Then, we finetune it for 500 epochs on the target kernels. We use 50 data points per kernel as the training set during fine-tune, and the other data points are the test set. We report the test MSE during the epoch which has the lowest training MSE. During fine-tuning, we use the same random seed as used in that model’s pretraining. We run offline evaluation five times. For online evaluation, since HLS takes a long time, we only use the three models trained with random seeds 1, 2, and 3 to run HLS. In offline evaluation, we calculate the sum of the MSE of the five prediction targets on the target kernels’ test set. In online evaluation, from the designs that are predicted to be valid and under resource constraints, we pick the top 10 designs with the lowest predicted latencies (highest predicted normalized latencies) to run HLS. When running HLS, we set the timeout as 3 hours. If the HLS does not finish within 3 hours, it is very likely to be unsynthesizable, so we will terminate the HLS and treat it as an invalid design. We report the best design from the selected top-10 and the training dataset of the target kernels, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We calculate the mean speedup of the three repeated HLS experiments for each target kernel. Then we calculate the geometric mean speedup for all target kernels.
@@
Baseline methods. We have three baseline models: HARP (Sohrabizadeh et al. 2023), HARP+MAML (Finn, Abbeel, and Levine 2017), and ProgSG (Qin et al. 2024). For our models, we follow the same hyper-parameter setting and training strategy as HARP, except for the MoE’s unique hyper-parameters. We will specify the hyper-parameters in a later section.
  • HARP (Sohrabizadeh et al. 2023): we use its original settings and official codes (https://github.com/UCLA-VAST/HARP).
  • HARP+MAML: inspired from the paper (Bai et al. 2022), we use the meta-learning method, MAML (Finn, Abbeel, and Levine 2017), to help find a good initialization for the HARP model’s parameters based on the source kernels. Then we use the normal fine-tuning strategy on the target kernels. We use the implementation of the “learn2learn” Python package. We use the same hyper-parameter setting as HARP. For the MAML part, we tune its hyper-parameters and select the hyper-parameters as follow: we randomly sample 32 kernels with replacement per epoch to train, and we use 80% of data as the training set.
  • ProgSG (Qin et al. 2024): it uses both text embedding of the source code and the ProGraML graph as the input. We use its official codes (https://github.com/zongyueqin/progsg).